OpenStudy (anonymous):

please help me understand how to solve this and these types of problems. 15 times the square of a non-zero number is equal to 15 times the number what is the number

OpenStudy (anonymous):

First, you have to know how to translate what you're being told to math language (algebra) and that will be, "15 times the square of a non-zero number" means 15x^2 for x not being 0. "is equal to 15 times the number" means it's an equation and the right side have 15 times x, so \[15*x^2=15*x\], x different from zero, so divide 15 and you'll have?

OpenStudy (anonymous):

3 cuz i divided 15 by 5

OpenStudy (anonymous):

why would you divide 15 by 5?

OpenStudy (anonymous):

cuz 3 and 5 are the only ones 15 can go in to evenly

OpenStudy (anonymous):

but look at your equation, you have 15x^2=15x then to eliminate constant, you must divide by...

OpenStudy (anonymous):

2

OpenStudy (anonymous):

okay, no you've \[15x^2=15x\] then divide by 15 \[(15/15) x^2=(15/15) x\] then, like (15/15)=1 \[x^2=x\] and the only number that is equal to his square is...

OpenStudy (anonymous):

1

OpenStudy (anonymous):

correct, did you understand the proccess?

OpenStudy (anonymous):

yes